# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@ 6. Numerical Experiments
In this experiment we consider RS-HB, HB-fastest, GD-pop, NAG-beta-opt, and TMM (using the parameters provided in table˜1 for each method). On the left plot of fig.˜5, we consider unbiased stochastic noise setting, where batching is used with batch size $b<N$ to construct an unbiased estimator of the gradient, i.e., the gradients are estimated with $\tilde{\nabla}f(y_{k})\coloneqq\frac{p}{b}\sum_{i=1}^{b}g(a_{I_{i}}^{\top}y_{k}+b_{I_{i}})+\mu y_{k}$, using $I_{1},...,I_{b}$ that are sampled with replacement from $\{1,...,N\}$ independently at step $k$. This type of estimation is common in the context of stochastic gradient descent and momentum methods [52, 113, 114]. For each method, we run $n=100$ sample paths, i.e., for every run $\ell$, $\{x_{j}^{(\ell)}\}_{j=0}^{K}$ is a simulated trajectory and each $\ell\in\{1,2,\dots,n\}$ corresponds to an independent run where we execute the algorithm in consideration with the iteration budget of $K=1000$ iterations. We estimate the sensitive cost sequence $\{R_{k}(\theta)\}_{k}$ numerically by considering the sample averages, i.e.,
@@
where $\hat{\sigma}^{2}$ is an estimate of the variance proxy $\sigma^{2}$ using the empirical variance of stochastic gradients.At iteration $k$, we compute the empirical variance, $\text{emp var}_{k}$, using 100 samples of the gradient evaluated at $y_{k}$, each of batch size 64. The final variance proxy estimate is $\hat{\sigma}^{2}\coloneqq\frac{1}{4}\max_{1\leq k\leq K}\text{emp var}_{k}$ –the scaling by $\frac{1}{4}$ is from the fact for a $\sigma^{2}$-subGaussian random variable, $X$, $\text{Var}[X]\leq 4\sigma^{2}$. This can be seen from the intermediary moment calculation in lemma E.1 by taking $\tilde{p}=2$. On the left plot of fig.˜5, we illustrate $\hat{R}_{k}(\theta)$ for $\theta\in\{0.001,1000\}$ in a logarithmic scale. We observe that GD-pop begins with a higher risk-sensitive cost than its momentum-based competitors, but after 1500 iterations this cost begins to fall below the competitors’ costs. This phenomenon reflects the trade-off between convergence rate and risk, as discussed in section˜4. While GD-pop converges more slowly, it can exhibit greater robustness to noise, since the impact of noise accumulates more gradually over iterations than the others. Furthermore, both HB-fastest and its robustly-stable variant (RS-HB) converge to higher risk levels than TMM and NAG-beta-opt, with NAG-beta-opt achieving the lowest risk among the momentum-based methods we compared. This behavior is consistent with the patterns observed in the quadratic case, shown in fig.˜3.
On the right plot of fig.˜5, we retain the same experimental setup as in the left plot except for one key difference: we introduce a bias term into the stochastic gradients. Specifically, at each iteration $k$, we add an adversarial (bias) component to the previously unbiased stochastic gradients. To construct this bias, for every iteration $k$, we sample 50 points from a $\delta$-ball centered at $\nabla f(y_{k})$ with $\delta=2.5$. This value of $\delta$ is chosen to clearly illustrate the contrast between the non-adversarial and adversarial scenarios. Among the sampled points, we select the one that maximizes the suboptimality $f(x_{k+1})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$. This form of adversarial perturbation is inspired by techniques from adversarial learning [47]. We plot the risk for both $\theta\in\{0.001,1000\}$. On the right plot, we observe the hierarchy in the performance of the algorithms; moreover, GD notably settles below its momentum-based competitors earlier (at around 750 iterations) than in the case with no adversarial noise.
@@ 7. Conclusion
We analyzed the trade-off between convergence rate and robustness to gradient errors in generalized momentum methods (GMMs), encompassing Nesterov’s accelerated gradient, heavy-ball, and gradient descent methods. Using the concept of risk sensitivity index (RSI) from robust control, we characterized robustness under stochastic, potentially biased, additive gradient errors. For quadratic objectives with isotropic Gaussian noise, we derived closed-form RSI expressions via Riccati equations and identified a Pareto frontier between convergence rate and the robustness. We further established a large-deviation principle that connects the rate function characterizing the associated large-deviation principle to RSI and the $H_{\infty}$ norm; thus, bridging robustness to stochastic biased noise with robustness to deterministic worst-case errors. Extending our results to strongly convex smooth objectives, beyond quadratics, we provided finite-time RSI bounds under biased sub-Gaussian errors, yielding high-probability and non-asymptotic guarantees, and highlighted similar robustness–convergence trade-offs. To our knowledge, this is the first risk-sensitive, non-asymptotic analysis of GMMs with biased gradients. We provided numerical experiments on robust regression to illustrate our results.
@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
[…]
@@
Since $f$ is a quadratic, the GMM system (4.5)–(4.7) is a linear dynamic system described by the system matrices $(A_{Q},B,T)$. For linear systems, it is well-known that the $H_{\infty}$ norm admits the frequency domain characterization
@@
(see e.g. [57]); where $G$ is the transfer matrix of the system $(A_{Q},B,T)$ defined as $G(z)\coloneqq T(z\mathsf{I}_{2d}-A_{Q})^{-1}B$ for $z\in\mathbb{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\setminus}\mbox{\text{Spec}}(A_{Q})$ with $\mbox{\text{Spec}}(A_{Q}):=\bigcup_{i}\{\lambda_{i}\}$ denoting the set of eigenvalues of $A_{Q}$. Thus, we can conclude that $\|G\|_{\infty}=H_{\infty}$ and $H_{\infty}$ is given by (3.16).
@@
$\mathrm{Ent}(G,\gamma)=\begin{cases}-\gamma^{2}\log\det(\mathsf{I}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}-\gamma^{-2}B^{\top}XB)&\mbox{if}\quad\|G\|_{\infty}<\gamma,\\ \infty&\mbox{if}\quad\|G\|_{\infty}\geq\gamma,\end{cases}$ (B.7)
where $X$ is the stabilizing solution to the discrete-time Riccati equation (4.8), i.e., $X$ solves (4.8) and satisfies $\rho(A_{Q}+\gamma^{-2}BB^{\top}X)<1$ (see e.g. [63], [92, Remark 4.11], [93, Lemma 3.4]). Applying this formula to (B.5) and using $\|G\|_{\infty}=H_{\infty}$, we conclude. ∎
@@
(see [109, Corollary 2.1, Prop. 3.7], [93, Lemma 3.4]) where we also know $\|G^{(\lambda_{i})}\|_{\infty}$ equals the $H_{\infty}$ norm of the corresponding system $(\tilde{A}^{(\lambda_{i})},\tilde{B},T^{(\lambda_{i})})$ (see (B.1)). This system subject to a given deterministic noise input $\{\tilde{\delta}_{k}\}_{k\geq 0}$ admits the state-space representation
@@
where in the second equality we used the fact that $Y$ is orthogonal, in the third equality we used $\overline{X}=Y\hat{X}Y^{\top}$ and in the last equality we used the fact that $Y^{\top}BB^{\top}Y=\hat{B}\hat{B}^{\top}$ and (B.14). We conclude from (B.22) that $\overline{X}$ is a stabilizing solution to (4.8). Finally, if the stabilizing solution exists then it is unique, see [115], [93, Lemma 3.4]. This completes the proof. ∎
@@
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{\theta}\log\det(\mathsf{I}_{d}-\frac{\theta}{d}B^{\top}Y\underset{i=1,...,n}{{\mbox{Diag}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{X}^{(\lambda_{i})}})Y^{\top}B)&\mbox{if}\quad\sqrt{\theta}H_{\infty}<\sqrt{d},\\ \infty&\mbox{if}\quad\sqrt{\theta}H_{\infty}\geq\sqrt{d},\end{cases}$ (B.23)
@@
Appendix C Proof of proposition˜4.3 and corollary˜4.4
C.1. Proof of proposition˜4.3
@@
It is known that the function $\tilde{\Lambda}_{K}(\lambda)$ is convex in $\lambda$ in its domain for every $K$.Indeed, for any $\bar{\lambda}_{1},\bar{\lambda}_{2}\in\mbox{dom}(\Lambda_{k}):=\{\lambda:\Lambda_{k}(\lambda)<\infty\}$, and $\omega\in[0,1]$, we have $\tilde{\Lambda}_{K}(\omega\bar{\lambda}_{1}+(1-\omega)\bar{\lambda}_{2})=\log\mathbb{E}[e^{\omega\bar{\lambda}_{1}\tilde{S}_{K}}e^{(1-\omega)\bar{\lambda}_{2}\tilde{S}_{K}}]\leq\log\big{(}\mathbb{E}[e^{\bar{\lambda}_{1}\tilde{S}_{K}}]^{\omega}\mathbb{E}[e^{\bar{\lambda}_{2}\tilde{S}_{K}}]^{1-\omega}\big{)}=\omega\tilde{\Lambda}_{K}(\bar{\lambda}_{1})+(1-\omega)\tilde{\Lambda}_{K}(\bar{\lambda}_{2})$ where the inequality is due to Hölder’s inequality. Thus, $\tilde{\Lambda}_{K}(\lambda)$ is convex in $\lambda$. Recall that the risk-sensitive index is given by: $R(\theta)=\lim_{K\to\infty}\frac{2\sigma^{2}}{\theta(K+1)}\log\mathbb{E}\left[e^{\frac{\theta}{2\sigma^{2}}S_{K}}\right]$. Therefore, taking $\tilde{S}_{K}\coloneqq\frac{1}{K+1}S_{K}$ and $\lambda=\frac{\theta}{2\sigma^{2}}$, we observe that the limit,
@@
where we used (C.2) and (C.3). Furthermore, $\tilde{\Lambda}$ is a convex function of $\lambda$ as it is the pointwise limit of convex functions $\frac{1}{K+1}\tilde{\Lambda}_{K}(\lambda)$. Gärtner-Ellis theorem [64, Section 2] says that if
@@
$I(s):=\sup_{\lambda\in\text{dom}(\Lambda)}[\lambda s-\tilde{\Lambda}(\lambda)]=\sup_{-\infty<\theta<\frac{d}{H_{\infty}^{2}}}\left[\frac{\theta}{2\sigma^{2}}\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}-R(\theta)\big{)}\right],$
which implies (4.17) as desired. In the rest of the proof, we will argue that these conditions hold. By the assumption $\rho(A_{Q})<1$, theorem˜3.4 applies and $H_{\infty}<\infty$. Using this fact and (C.4), we conclude that condition (I) and condition (i) holds. From the representation (C.1), we observe that $R(\theta)$ is continuous and differentiable in its domain; therefore, using (C.3), we conclude that $\tilde{\Lambda}$ is continuous and differentiable on its domain. Hence, conditions (II) and (ii) both also hold. It suffices to argue that condition (iii) hold. Note that $\lambda_{*}=\frac{d}{2\sigma^{2}H_{\infty}^{2}}$ is the only boundary point for the domain of $\tilde{\Lambda}$ and
@@
Therefore, combining (C.5), (C.9) and (C.8), we obtain $\tilde{\Lambda}^{\prime}(\lambda_{n})=R(\theta_{n})+2\sigma^{2}\lambda R^{\prime}(\theta_{n})\to\infty,$ and this proves (iii) as desired and the LDP follows from the Gärtner-Ellis Theorem. To prove (4.19), note that by the Gaussian noise assumption, the density of $S_{K}/(K+1)$ is positive and continuous at any point $s\in(0,\infty)$; therefore, $I(s)<+\infty$ for any $s>0$ (see [64, Ex. 2.2.39]. Thus, the rate function $I$, as the dual of $\tilde{\Lambda}$, is continuous on its domain which include $(0,\infty)$, and we conclude that for the set $E=[t,\infty]$ with $t>0$,
@@
C.2. Proof of corollary˜4.4
For momentum methods, [57] shows that the transfer matrix $G(e^{i\omega})$ admits the explicit representation
@@
Then, exploiting the diagonal structure of $G(e^{i\omega})G^{*}(e^{i\omega})$, results in the equalities
@@
$\int_{-\pi}^{\pi}\operatorname{trace}\left(G(e^{i\omega})G^{*}(e^{i\omega})\right)d\omega=\displaystyle\sum_{j=1}^{d}\displaystyle\int_{-\pi}^{\pi}h_{\omega}(\lambda_{i}),d\omega,$
showing that (4.18) is equivalent to (4.20) and completes the proof.
@@
From (D.2) and the definition of the risk-sensitive index, we can infer thatWe use the convention that the right-hand side of (D.5) is $\infty$, when $\frac{\bar{\theta}}{J_{p,q}}\geq\bar{\theta}^{ub}$; in which case the bound (D.5) is trivial.
@@
for suitably chosen non-negative scalars $a_{k,K},b_{k,K}$ and $c_{k,K}$ (that may depend on the choice of $\bar{\theta}$) where $\xi^{+}$ and $\xi$ are as in (D.3). Clearly, such a bound holds for $k=K$ with $a_{K,K}=J_{p,q},b_{K,K}=1-J_{p,q}$ and $c_{K,K}(\bar{\theta})=0$. Next, we consider $k<K$, and introduce the next GMM iterate starting from $\xi_{k}=\xi^{+}$:

### v2

@@ 6. Numerical Experiments
In this experiment we consider RS-HB, HB-fastest, GD-pop, NAG-beta-opt, and TMM (using the parameters provided in table˜1 for each method). On the left plot of fig.˜5, we consider unbiased stochastic noise setting, where batching is used with batch size $b<p$ to construct an unbiased estimator of the gradient, i.e., the gradients are estimated with $\tilde{\nabla}f(y_{k})\coloneqq\frac{p}{b}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}=1}^{b}g_{\lambda}(a_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i_{t}}}^{\top}y_{k}+b_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i_{t}}})+\mu y_{k}$, using $i_{1},...,i_{b}$ that are sampled with replacement from $\{1,...,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}\}$ independently at step $k$. This type of estimation is common in the context of stochastic gradient descent and momentum methods [52, 116, 117]. For each method, we run $n=100$ sample paths, i.e., for every run $\ell$, $\{x_{j}^{(\ell)}\}_{j=0}^{K}$ is a simulated trajectory and each $\ell\in\{1,2,\dots,n\}$ corresponds to an independent run where we execute the algorithm in consideration with the iteration budget of $K=1000$ iterations. We estimate the sensitive cost sequence $\{R_{k}(\theta)\}_{k}$ numerically by considering the sample averages, i.e.,
@@
where $\hat{\sigma}^{2}$ is an estimate of the variance proxy $\sigma^{2}$ using the empirical variance of stochastic gradients. [^5] At iteration $k$, we compute the empirical variance, $\text{emp var}_{k}$, using 100 samples of the gradient evaluated at $y_{k}$, each of batch size 64. The final variance proxy estimate is $\hat{\sigma}^{2}\coloneqq\frac{1}{4}\max_{1\leq k\leq K}\text{emp var}_{k}$ –the scaling by $\frac{1}{4}$ is from the fact for a $\sigma^{2}$-subGaussian random variable, $X$, $\text{Var}[X]\leq 4\sigma^{2}$. This can be seen from the intermediary moment calculation in lemma E.1 by taking $\tilde{p}=2$. On the left plot of fig.˜5, we illustrate $\hat{R}_{k}(\theta)$ for $\theta\in\{0.001,1000\}$ in a logarithmic scale. We observe that GD-pop begins with a higher risk-sensitive cost than its momentum-based competitors, but after 1500 iterations this cost begins to fall below the competitors’ costs. This phenomenon reflects the trade-off between convergence rate and risk, as discussed in section˜4. While GD-pop converges more slowly, it can exhibit greater robustness to noise, since the impact of noise accumulates more gradually over iterations than the others. Furthermore, both HB-fastest and its robustly-stable variant (RS-HB) converge to higher risk levels than TMM and NAG-beta-opt, with NAG-beta-opt achieving the lowest risk among the momentum-based methods we compared. This behavior is consistent with the patterns observed in the quadratic case, shown in fig.˜3.
On the right plot of fig.˜5, we retain the same experimental setup as in the left plot except for one key difference: we introduce a bias term into the stochastic gradients. Specifically, at each iteration $k$, we add an adversarial (bias) component to the previously unbiased stochastic gradients. To construct this bias, for every iteration $k$, we sample 50 points from a $\delta$-ball centered at $\nabla f(y_{k})$ with $\delta=2.5$. This value of $\delta$ is chosen to clearly illustrate the contrast between the non-adversarial and adversarial noise scenarios. Among the sampled points, we select the one that maximizes the suboptimality $f(x_{k+1})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$. This form of adversarial perturbation is inspired by techniques from adversarial learning [47]. We plot the risk for both $\theta\in\{0.001,1000\}$. On the right plot, we observe the hierarchy in the performance of the algorithms; moreover, GD notably settles below its momentum-based competitors earlier (at around 750 iterations) than in the case with no adversarial noise.
@@ 7. Conclusion
We analyzed the trade-off between convergence rate and robustness to gradient errors in generalized momentum methods (GMMs), encompassing Nesterov’s accelerated gradient, heavy-ball, and gradient descent methods. Using the concept of risk sensitivity index (RSI) from robust control, we characterized robustness under stochastic, potentially biased, additive gradient errors. For quadratic objectives with isotropic Gaussian noise, we derived closed-form RSI expressions via Riccati equations and identified a Pareto frontier between convergence rate and the robustness. We further established a large-deviation principle that connects the rate function characterizing the associated large-deviation principle to RSI and the $H_{\infty}$-norm, thus bridging robustness to stochastic biased noise with robustness to deterministic worst-case errors. Extending our results to strongly convex smooth objectives, beyond quadratics, we provided finite-time RSI bounds under biased sub-Gaussian errors, yielding high-probability and non-asymptotic guarantees, and highlighted similar robustness–convergence trade-offs. To our knowledge, this is the first risk-sensitive, non-asymptotic analysis of GMMs with biased gradients. We provided numerical experiments on robust regression to illustrate our results.
@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
$\small H_{\infty}\leq\overline{H}_{\infty}=\begin{cases}\frac{1}{\sqrt{2\mu}}&\mbox{if}\qquad 0<\alpha\leq\frac{1}{L},\\ \frac{1}{\sqrt{2\mu}}\frac{\alpha L}{2-\alpha L}&\mbox{if}\qquad\frac{1}{L}<\alpha\leq\frac{2}{L+\sqrt{L\mu}},\\ \frac{1}{\sqrt{2\mu}}\sqrt{\kappa}&\mbox{if}\qquad\frac{2}{L+\sqrt{L\mu}}<\alpha\leq\frac{2}{L+\mu},\\ \frac{\alpha\sqrt{L}}{\sqrt{2}(2-\alpha L)}&\mbox{if}\qquad\frac{2}{L+\mu}<\alpha<\frac{2}{L}.\end{cases}$
@@
Since $f$ is a quadratic, the GMM system (4.5)–(4.7) is a linear dynamic system described by the system matrices $(A_{Q},B,T)$. For linear systems, it is well-known that the $H_{\infty}$-norm admits the frequency domain characterization
@@
(see e.g., [57]); where $G$ is the transfer matrix of the system $(A_{Q},B,T)$ defined as $G(z)\coloneqq T(z\mathsf{I}_{2d}-A_{Q})^{-1}B$ for $z\in\mathbb{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\setminus}\mbox{\text{Spec}}(A_{Q})$ with $\mbox{\text{Spec}}(A_{Q}):=\bigcup_{i}\{\lambda_{i}\}$ denoting the set of eigenvalues of $A_{Q}$. Thus, we can conclude that $\|G\|_{\infty}=H_{\infty}$ and $H_{\infty}$ is given by (3.16).
@@
$\mathrm{Ent}(G,\gamma)=\begin{cases}-\gamma^{2}\log\det(\mathsf{I}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}-\gamma^{-2}B^{\top}XB)&\mbox{if}\qquad\|G\|_{\infty}<\gamma,\\ \infty&\mbox{if}\qquad\|G\|_{\infty}\geq\gamma,\end{cases}$ (B.7)
where $X$ is the stabilizing solution to the discrete-time Riccati equation (4.8), i.e., $X$ solves (4.8) and satisfies $\rho(A_{Q}+\gamma^{-2}BB^{\top}X)<1$ (see e.g., [64], [93, Remark 4.11], [94, Lemma 3.4]). Applying this formula to (B.5) and using $\|G\|_{\infty}=H_{\infty}$, we conclude. ∎
@@
(see [112, Corollary 2.1, Prop. 3.7], [94, Lemma 3.4]) where we also know $\|G^{(\lambda_{i})}\|_{\infty}$ equals the $H_{\infty}$-norm of the corresponding system $(\tilde{A}^{(\lambda_{i})},\tilde{B},T^{(\lambda_{i})})$ (see (B.1)). This system subject to a given deterministic noise input $\{\tilde{\delta}_{k}\}_{k\geq 0}$ admits the state-space representation
@@
where in the second equality we used the fact that $Y$ is orthogonal, in the third equality we used $\overline{X}=Y\hat{X}Y^{\top}$ and in the last equality we used the fact that $Y^{\top}BB^{\top}Y=\hat{B}\hat{B}^{\top}$ and (B.14). We conclude from (B.22) that $\overline{X}$ is a stabilizing solution to (4.8). Finally, if the stabilizing solution exists then it is unique, see [118], [94, Lemma 3.4]. This completes the proof. ∎
@@
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{\theta}\log\det(\mathsf{I}_{d}-\frac{\theta}{d}B^{\top}Y\underset{i=1,...,d}{{\mbox{Diag}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{X}^{(\lambda_{i})}})Y^{\top}B)&\mbox{if}\quad\sqrt{\theta}H_{\infty}<\sqrt{d},\\ \infty&\mbox{if}\quad\sqrt{\theta}H_{\infty}\geq\sqrt{d},\end{cases}$ (B.23)
@@
Appendix C Proofs of proposition˜4.4 and corollary˜4.6
C.1. Proof of proposition˜4.4
@@
It is known that the function $\tilde{\Lambda}_{K}(\lambda)$ is convex in $\lambda$ in its domain for every KK. [^9] Indeed, for any $\bar{\lambda}_{1},\bar{\lambda}_{2}\in\mbox{dom}(\Lambda_{k}):=\{\lambda:\Lambda_{k}(\lambda)<\infty\}$, and $\omega\in[0,1]$, we have $\tilde{\Lambda}_{K}(\omega\bar{\lambda}_{1}+(1-\omega)\bar{\lambda}_{2})=\log\mathbb{E}[e^{\omega\bar{\lambda}_{1}\tilde{S}_{K}}e^{(1-\omega)\bar{\lambda}_{2}\tilde{S}_{K}}]\leq\log\big(\mathbb{E}[e^{\bar{\lambda}_{1}\tilde{S}_{K}}]^{\omega}\mathbb{E}[e^{\bar{\lambda}_{2}\tilde{S}_{K}}]^{1-\omega}\big)=\omega\tilde{\Lambda}_{K}(\bar{\lambda}_{1})+(1-\omega)\tilde{\Lambda}_{K}(\bar{\lambda}_{2})$ where the inequality is due to Hölder’s inequality. Thus, $\tilde{\Lambda}_{K}(\lambda)$ is convex in $\lambda$. Recall that the risk-sensitive index is given by: $R(\theta)=\lim_{K\to\infty}\frac{2\sigma^{2}}{\theta(K+1)}\log\mathbb{E}\left[e^{\frac{\theta}{2\sigma^{2}}S_{K}}\right]$. Therefore, taking $\tilde{S}_{K}\coloneqq\frac{1}{K+1}S_{K}$ and $\lambda=\frac{\theta}{2\sigma^{2}}$, we observe that the limit,
@@
where we used (C.2) and (C.3). Furthermore, $\tilde{\Lambda}$ is a convex function of $\lambda$ as it is the pointwise limit of convex functions $\frac{1}{K+1}\tilde{\Lambda}_{K}(\lambda)$. Gärtner-Ellis theorem [65, Section 2] says that if
@@
$I(s):=\sup_{\lambda\in\text{dom}(\tilde{\Lambda})}[\lambda s-\tilde{\Lambda}(\lambda)]=\sup_{-\infty<\theta<\frac{d}{H_{\infty}^{2}}}\left[\frac{\theta}{2\sigma^{2}}\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}-R(\theta)\big)\right],$
which implies (4.17) as desired. In the rest of the proof, we will argue that these conditions hold. By the assumption $\rho(A_{Q})<1$, theorem˜3.4 applies and $H_{\infty}<\infty$. Using this fact and (C.4), we conclude that condition (I) and condition (i) holds. From the representation (C.1), we observe that $R(\theta)$ is continuous and differentiable in its domain; therefore, using (C.3), we conclude that $\tilde{\Lambda}$ is continuous and differentiable on its domain. Hence, conditions (II) and (ii) both also hold. It suffices to argue that condition (iii) holds. Note that $\lambda_{*}=\frac{d}{2\sigma^{2}H_{\infty}^{2}}$ is the only boundary point for the domain of $\tilde{\Lambda}$ and
@@
Therefore, combining (C.5), (C.9) and (C.8), we obtain $\tilde{\Lambda}^{\prime}(\lambda_{n})=R(\theta_{n})+2\sigma^{2}\lambda R^{\prime}(\theta_{n})\to\infty,$ and this proves (iii) as desired and the LDP follows from the Gärtner-Ellis Theorem. To prove (4.21), note that by the Gaussian noise assumption, the density of $S_{K}/(K+1)$ is positive and continuous at any point $s\in(0,\infty)$; therefore, $I(s)<+\infty$ for any $s>0$ (see [65, Ex. 2.2.39]. Thus, the rate function $I$, as the dual of $\tilde{\Lambda}$, is continuous on its domain which includes $(0,\infty)$, and we conclude that for the set $E=[t,\infty]$ with $t>0$,
@@
C.2. Proof of corollary˜4.6
For momentum methods, [57, eqn. (59)] shows that the transfer matrix $G(e^{i\omega})$ admits the explicit representation
@@
Then, by exploiting the diagonal structure of $G(e^{i\omega})G^{*}(e^{i\omega})$, we obtain the equalities
@@
$\int_{-\pi}^{\pi}\operatorname{trace}\left(G(e^{i\omega})G^{*}(e^{i\omega})\right)d\omega=\displaystyle\sum_{j=1}^{d}\displaystyle\int_{-\pi}^{\pi}h_{\omega}(\lambda_{i})d\omega,$
showing that (4.18) is equivalent to (4.22). This completes the proof.
@@
From (D.2) and the definition of the risk-sensitive cost, we can infer that [^10] We use the convention that the right-hand side of (D.5) is $\infty$, when $\frac{\bar{\theta}}{J_{p,q}}\geq\bar{\theta}^{ub}$; in which case the bound (D.5) is trivial.
@@
for suitably chosen scalars $a_{k,K},b_{k,K}$ and $c_{k,K}$ (that may depend on the choice of $\bar{\theta}$) where $\xi^{+}$ and $\xi$ are as in (D.3). Clearly, such a bound holds for $k=K$ with $a_{K,K}=J_{p,q},b_{K,K}=1-J_{p,q}$ and $c_{K,K}(\bar{\theta})=0$. Next, we consider $k<K$, and introduce the next GMM iterate starting from $\xi_{k}=\xi^{+}$: